OpenStudy (anonymous):

Verify the identity. cot(x-pi/2) = -tan x

terenzreignz (terenzreignz):

\[\Large \cot\left(x - \frac{\pi}{2}\right)=\frac{\cos\left(x - \frac{\pi}{2}\right)}{\sin\left(x - \frac{\pi}{2}\right)}\]

terenzreignz (terenzreignz):

@Jdina97 make your presence felt :3

OpenStudy (anonymous):

im here

OpenStudy (anonymous):

How do you break down the inside. What's in the parentheses

terenzreignz (terenzreignz):

Okay, so you know that \[\Large \cos(x) = \cos(-x) \]right? :) also, \[\Large \sin(x) =-\sin(-x)\] These will make things easier :)

OpenStudy (anonymous):

yup

terenzreignz (terenzreignz):

So the numerator can be written as: \[\Large \cot\left(x - \frac{\pi}{2}\right)=\frac{\cos\left(\frac{\pi}2 - x\right)}{-\sin\left(\frac{\pi}2 - x\right)}\]

OpenStudy (anonymous):

but doesn't the Sin(x) =-sin(x) not -sin(-x)?

terenzreignz (terenzreignz):

Oh no :) sin(-x) = -sin(x) So it only follows that sin(x) = -sin(-x)

terenzreignz (terenzreignz):

You want me to prove it? :3

OpenStudy (anonymous):

Oh okay my bad i saw it wrong

terenzreignz (terenzreignz):

all righty then

OpenStudy (anonymous):

I misunderstood what my Formula sheet said:P

terenzreignz (terenzreignz):

Okay, so another thing that's useful is the fact that \[\Large \sin\left(\frac{\pi}{2} - x\right)= \cos(x) \]\[\Large \cos\left(\frac{\pi}{2} - x\right) = \sin(x)\]

OpenStudy (anonymous):

Hmm. How did you get that?

OpenStudy (anonymous):

Thank for answering by the way

terenzreignz (terenzreignz):

:) \[\Large \sin(\alpha-\beta) = \sin(\alpha)\cos(\beta) - \cos(\alpha)\sin(\beta) \]

terenzreignz (terenzreignz):

Now if we let \[\Large \alpha = \frac\pi 2\]\[\Large \beta = x\]

OpenStudy (anonymous):

Oh okay

terenzreignz (terenzreignz):

We get \[\Large \sin\left(\frac{\pi}2\right)\cos(x) -\cos\left(\frac{\pi}2\right)\sin(x)\]

terenzreignz (terenzreignz):

\[\Large \sin\left(\frac{\pi}2\right) = 1\]\[\Large \cos\left(\frac{\pi}2\right)=0\] Therefore... \[\Large = \cos(x) \] Understood ? :)

OpenStudy (anonymous):

Sort of

terenzreignz (terenzreignz):

Well, that's not good enough :) What part don't you get? :D

OpenStudy (anonymous):

I understood how you got sin(π2)cos(x)−cos(π2)sin(x)

OpenStudy (anonymous):

Oh wait. After that you go to the unit circle correct?

terenzreignz (terenzreignz):

Yes, if only to find out that \[\large \sin\left(\frac{\pi}2\right)= 1\]and\[\large \cos\left(\frac{\pi}2\right)=0\]

OpenStudy (anonymous):

Okay yea that makes sense now.

terenzreignz (terenzreignz):

so, similarly, we get \[\large \cos\left(\frac{\pi}2-x\right)=\cos\left(\frac{\pi}2\right)\cos(x)+\sin\left(\frac{\pi}2\right)\sin(x)=\sin(x)\]

terenzreignz (terenzreignz):

So, no more doubts that \[\sin\left(\frac{\pi}2-x\right)=\cos(x)\]and\[\cos\left(\frac{\pi}2-x\right)=\sin(x)\]

OpenStudy (anonymous):

Yup

terenzreignz (terenzreignz):

Okay, so back to this \[\Large \cot\left(x - \frac{\pi}{2}\right)=\frac{\cos\left(\frac{\pi}2 - x\right)}{-\sin\left(\frac{\pi}2 - x\right)}\]

OpenStudy (anonymous):

And then Sin/Cos = Tan and Since Sin is Negative it equals -tanx correct?

terenzreignz (terenzreignz):

Very observant :) And your observation is correct ^_^

OpenStudy (anonymous):

Thank you so much for your help :D

terenzreignz (terenzreignz):

\[\Large \cot\left(x - \frac{\pi}{2}\right)=\frac{\cos\left(\frac{\pi}2 - x\right)}{-\sin\left(\frac{\pi}2 - x\right)}=\frac{\sin(x)}{-\cos(x)}=-\tan(x)\]

OpenStudy (anonymous):

Now do you mind helping someone else :P i started helping them but i realized i was saying the wrong thing. They needed help with a "train station problem" from geometry but i forget how to do it

terenzreignz (terenzreignz):

I'll see... link/tag me :)